# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v2

@@ 3.2 Stepsize bounds
In particular, $\{V_{k}\}$ is nonincreasing if $\alpha\leq\bar{\alpha}$, $\beta\leq L_{{}_{\rm{p}}}^{-1}$ and $\gamma\leq\rho_{{}_{\rm{g}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
$V_{k}\leq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})-\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}+g(y^{k}).$ (3.28)
@@
$\begin{split}V_{k-1}&\geq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})+g(y^{k})\\ &\quad+\left\langle\nabla f(x^{k})-\nabla f(x^{k-1})-\frac{1}{\alpha}(x^{k}-x^{k-1}),x^{k-1}-y^{k}\right\rangle\\ &\quad+\left\langle\nabla h(x^{k-1})-\nabla h(x^{k}),y^{k}-x^{k}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.\end{split}$ (3.29)
@@
$\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$
$\quad+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}.$ (3.30)
@@
$\quad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}.$ (3.37)
@@
$\quad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}$
$=-\frac{c(\hat{\alpha})}{2\tau\hat{\alpha}}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$
$\quad+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$ (3.38)
@@
$\begin{split}c(\hat{\alpha})&\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\hat{\alpha}-(2-\tau)\\ &=\left(\hat{\alpha}-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)\left(2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}+(2-\tau)(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\right)\end{split},$ (3.39)
@@
$V_{k-1}-V_{k}$ $\geq-\frac{c(\hat{\alpha}(L))}{2\tau\hat{\alpha}(L)}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.41)
$\quad+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$
@@
and similar to the previous case, we get that $d(\bar{\alpha})>0$ and therefore $\hat{\alpha}<\bar{\alpha}$. This shows that our stepsize upper bound is always larger than that in (Bian and Zhang, 2021). The significant gap between the computed stepsizes is evident in Fig. 1.
@@
$V_{k-1}-V_{k}$ $\geq\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}{\left\|{y^{k}-y^{k+1}}\right\|}^{2},$
@@
The following provides a result similar to the previous theorem but for parameters $\tau$ in $(1,2)$.
@@
Suppose that Assumption 1.1 holds and $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$. Let $\sigma_{{}_{\rm{h}}}\in{\rm I}\!{\rm R}$ be such that $h-\frac{\sigma_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex. If $\{(x^{k},y^{k},z^{k})\}$ is generated by Algorithm 1 with $\tau\in(1,2)$ and $\alpha\in(0,\bar{\alpha}]$, where
@@
otherwise. In particular, $\{V_{k}\}$ is nonincreasing if $\alpha\leq\bar{\alpha}$, $\beta\leq L_{{}_{\rm{p}}}^{-1}$ and $\gamma\leq\rho_{{}_{\rm{g}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
We note that there indeed exists $\sigma_{{}_{\rm{h}}}\in{\rm I}\!{\rm R}$ such that $h-\frac{\sigma_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex; for instance, we can take $\sigma_{{}_{\rm{h}}}=-L_{{}_{\rm{h}}}$ by Lemma 2.1(a).
@@
where $\hat{c}(\alpha)\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)$. Since $\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ is the positive root of $\hat{c}(\alpha)$ (see (3.39)) and $L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}}\geq 0$ by Remark 2.2, it holds that $c\left(\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)=\frac{2(\tau-1)(L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}})}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\geq 0$. Since $c(0)=\tau-2<0$ and $\bar{\alpha}_{1}$ is the positive root of $c(\alpha)$, it follows that $\bar{\alpha}_{1}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. On the other hand, following the same arguments in Remark 3.6, it can be shown that $\tau/(2\eta^{*})\leq\bar{\alpha}_{1}$, so Eq. 3.45 holds as claimed.
@@
Suppose that Assumption 1.1 holds, $f$ is $\sigma_{{}_{\rm{f}}}$-strongly convex for some $\sigma_{f}>0$, and $\rho_{{}_{\rm{h}}}\in[0,L_{{}_{\rm{h}}}]$ such that $h+\frac{\rho_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex. Let $\tau\geq 2$ be such that
@@
In particular, $\{V_{k}\}$ is nonincreasing if $\frac{\tau\mu_{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}\leq\alpha\leq\frac{\tau\mu^{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$, $\beta\leq L_{{}_{\rm{p}}}^{-1}$ and $\gamma\leq\rho_{{}_{\rm{g}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
$\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}-\frac{(1-\mu)\sigma_{{}_{\rm{f}}}}{2}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$
@@
$\quad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}.$
@@
Suppose that $\tau=2$. Then Eqs. 3.48 and 3.49 are equivalent to having $\sigma_{{}_{\rm{f}}}>L_{{}_{\rm{h}}}+\rho_{{}_{\rm{h}}}$, and the roots of (3.50) are $\mu_{*}=0$ and $\mu^{*}=\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})(\sigma_{{}_{\rm{f}}}-L_{{}_{\rm{h}}}-\rho_{{}_{\rm{h}}})}{L_{{}_{\rm{h}}}(L_{{}_{\rm{f}}}^{2}-\sigma_{{}_{\rm{f}}}^{2})+\sigma_{{}_{\rm{f}}}L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$. Hence, $c(\alpha)$ is strictly negative for all $\alpha$ such that
@@ 3.3 Subsequential convergence
Suppose that Assumption 1.1 holds, $\Psi$ has bounded level sets, $\beta\in(0,L_{{}_{\rm{p}}}^{-1}]$, and $\gamma\in(0,\rho_{{}_{\rm{g}}}^{-1}]$. In addition, suppose that $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$, and $\alpha,\tau>0$ are chosen such that
@@
    $\tau\geq 2$ satisfies Eqs. 3.48 and 3.49, and $\alpha\in\left(\frac{\tau\mu_{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})},\frac{\tau\mu^{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}\right)\cap\left(0,\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)$.
@@
(3.23)${\geq}\Psi(y^{+})+\frac{1-\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2},$ (3.54)
@@
$V_{k-1}\geq\Psi(y^{k})+\frac{1-\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{2\alpha}{\left\|{y^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.$ (3.55)
@@
When $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}=0$, $x$ and $z$ play no role in the algorithm. We thus obtain the following inequality similar to (3.54):
$V_{\Lambda,\xi}(y,z,x)\geq\Psi(y^{+})+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$
Hence, if $\Psi$ has bounded level sets and $\beta<\frac{1}{L_{{}_{\rm{p}}}}$, we obtain the boundedness of $\{y^{k}\}$ and that ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$.
We will next show that accumulation points of the $x$ and $y$ sequence generated from Algorithm 1 are stationary points of $\Psi$. We say that a point $w^{*}$ is a stationary point of $\Psi$ (see (Wen et al., 2018, Definition 4.1)) if
$0\in\nabla f(w^{*})+\partial g(w^{*})+\nabla h(w^{*})+\partial p(w^{*}).$ (3.57)
Note that any local minimizer of $\Psi$ is a stationary point of $\Psi$ (see (Pham and Thi, 1997, Theorem 2(i))). To establish stationarity of accumulation points, we will prove that these points are fixed points of the defining operator of Algorithm 1. Observe that Algorithm 1 can be concisely written as fixed-point iterations of a certain map; in particular,
@@
where $T_{\Lambda}:{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}$ is given by
@@
and use it as a measure of stationarity from the view that $R(y,z)=0$ if and only if $(y,z)\in\operatorname*{Fix}(T_{\Lambda})$. This residual function reduces to that in (Lee and Wright, 2019) when $f\equiv 0$ and $p\equiv 0$, and to that in (Liu and Takeda, 2022) when $f\equiv 0$. A direct consequence of inequality (3.24) and Proposition 3.15 is the iteration complexity given in the following result.
@@
Suppose that the hypotheses of Proposition 3.15 hold, and let $(x^{*},y^{*},z^{*})$ be an accumulation point of $\{(x^{k},y^{k},z^{k})\}$. Then $x^{*}=y^{*}$ and $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$. Moreover, if either $\beta<L_{{}_{\rm{p}}}^{-1}$ or $\gamma<\rho_{{}_{\rm{g}}}^{-1}$, then
$\min_{k=1,2,\dots,N}R(y^{k},z^{k})^{2}\leq\frac{\omega}{N},$
where $\omega>0$ is given by
$\omega\coloneqq\begin{cases}\frac{2\beta(V_{0}-\Psi^{*})}{\min\left\{2\beta\zeta,1-\beta L_{{}_{\rm{p}}}\right\}}&\text{if}~{}\beta<L_{{}_{\rm{p}}}^{-1},\\ \frac{4\gamma(V_{0}-\Psi^{*})}{\min\left\{2\gamma\zeta,1-\gamma\rho_{{}_{\rm{g}}}\right\}}&\text{if}~{}\gamma<\rho_{{}_{\rm{g}}}^{-1},\end{cases}\quad\text{and}\quad\zeta\coloneqq\frac{-c(\alpha)}{2\tau\alpha(1+\alpha L_{{}_{\rm{f}}})^{2}}>0,$
and $\Psi^{*}\coloneqq\min_{w\in{\rm I}\!{\rm R}^{n}}\Psi(w)$.

### v4

@@ 3.2 Stepsize bounds
In particular, $\{V_{k}\}$ is nonincreasing if $\alpha\leq\bar{\alpha}$ and $\beta\leq\rho_{{}_{\rm{p}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
$V_{k}\leq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})+g(y^{k}).$ (3.20)
@@
$\begin{split}V_{k-1}&\geq Q_{\alpha(f+h)}(y^{k};x^{k})+p(y^{k})+g(y^{k})\\ &\quad+\left\langle\nabla f(x^{k})-\nabla f(x^{k-1})-\frac{1}{\alpha}(x^{k}-x^{k-1}),x^{k-1}-y^{k}\right\rangle\\ &\quad+\left\langle\nabla h(x^{k-1})-\nabla h(x^{k}),y^{k}-x^{k}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.\end{split}$ (3.21)
@@
$\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.22)
@@
$\quad+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.$ (3.29)
@@
$\quad+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$
$=-\frac{c(\hat{\alpha})}{2\tau\hat{\alpha}}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.30)
@@
$c(\hat{\alpha})$ $\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\hat{\alpha}-(2-\tau)$
$=\left(\hat{\alpha}-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)\left(2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}+(2-\tau)(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\right),$ (3.31)
@@
$V_{k-1}-V_{k}$ $\geq-\frac{c(\hat{\alpha}(L))}{2\tau\hat{\alpha}(L)}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.33)
@@
and similar to the previous case, we get that $d(\bar{\alpha})>0$ and therefore $\hat{\alpha}<\bar{\alpha}$. This shows that our stepsize upper bound is always larger than that in (Bian and Zhang, 2021). The significant gap between the computed stepsizes is evident in Fig. 1. As we will see in the experiment in Section 4.1.1, a larger $\alpha$ leads to faster convergence. Our improved upper bound for $\alpha$ therefore also provides numerical acceleration for DYS.
@@
$V_{k-1}-V_{k}$ $\geq\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2},$
@@
The following provides a result similar to the previous theorem but for parameters $\tau$ in $(1,2)$. Note that the parameter $\sigma_{{}_{\rm{h}}}$ in the theorem below exists; for instance, we can take $\sigma_{{}_{\rm{h}}}=-L_{{}_{\rm{h}}}$ by Lemma 2.1(a).
@@
Suppose that Assumption 1.1 holds and $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$. Let $\sigma_{{}_{\rm{h}}}\in{\rm I}\!{\rm R}$ be such that $h-\frac{\sigma_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.444 We note that there indeed exists $\sigma_{{}_{\rm{h}}}\in{\rm I}\!{\rm R}$ such that $h-\frac{\sigma_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex; for instance, we can take $\sigma_{{}_{\rm{h}}}=-L_{{}_{\rm{h}}}$ by Lemma 2.1(a). If $\{(x^{k},y^{k},z^{k})\}$ is generated by Algorithm 1 with $\tau\in(1,2)$ and $\alpha\in(0,\bar{\alpha}]$, where
@@
otherwise. In particular, $\{V_{k}\}$ is nonincreasing if $\alpha\leq\bar{\alpha}$ and $\beta\leq\rho_{{}_{\rm{p}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
where $\hat{c}(\alpha)\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)$. Since $\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ is the positive root of $\hat{c}(\alpha)$ (see (3.31)) and $L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}}\geq 0$ by Eq. 2.5, it holds that $c\left(\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)=\frac{2(\tau-1)(L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}})}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\geq 0$. Since $c(0)=\tau-2<0$ and $\bar{\alpha}_{1}$ is the positive root of $c(\alpha)$, it follows that $\bar{\alpha}_{1}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. On the other hand, following the same arguments in Remark 3.6, it can be shown that $\tau/(2\eta^{*})\leq\bar{\alpha}_{1}$, so Eq. 3.37 holds as claimed.
@@
Suppose that Assumption 1.1 holds and $\rho_{{}_{\rm{h}}}\in[0,L_{{}_{\rm{h}}}]$ such that $h+\frac{\rho_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex. [^5] By Assumption 1.1(a) and Eq. 2.5, we see that there indeed exists $\rho_{{}_{\rm{h}}}\geq 0$ such that $h+\frac{\rho_{{}_{\rm{h}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex. In addition, suppose that $f$ is $\sigma_{{}_{\rm{f}}}$-strongly convex for some $\sigma_{{}_{\rm{f}}}>0$. Let $\tau\geq 2$ be such that
@@
In particular, $\{V_{k}\}$ is nonincreasing if $\alpha\leq\bar{\alpha}$ and $\beta\leq\rho_{{}_{\rm{p}}}^{-1}$, and strictly decreasing if at least one holds with strict inequality.
@@
$\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}-\frac{(1-\mu)\sigma_{{}_{\rm{f}}}}{2}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$
@@
$\quad+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.44)
@@
Suppose that $\tau=2$. Then Eqs. 3.40 and 3.41 are equivalent to having $\sigma_{{}_{\rm{f}}}>L_{{}_{\rm{h}}}+\rho_{{}_{\rm{h}}}$, and the roots of (3.42) are $\mu_{*}=0$ and $\mu^{*}=\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})(\sigma_{{}_{\rm{f}}}-L_{{}_{\rm{h}}}-\rho_{{}_{\rm{h}}})}{L_{{}_{\rm{h}}}(L_{{}_{\rm{f}}}^{2}-\sigma_{{}_{\rm{f}}}^{2})+\sigma_{{}_{\rm{f}}}L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$. Hence, $c(\alpha)$ is strictly negative for all $\alpha$ such that
@@ 3.3 Subsequential convergence
Suppose that Assumption 1.1 holds, $\Psi$ has bounded level sets and $\beta\in(0,\rho_{{}_{\rm{p}}}^{-1}]$. In addition, suppose that $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$, and $\alpha,\tau>0$ are chosen such that
@@
    $\tau\geq 2$ satisfies Eqs. 3.40 and 3.41, and $\alpha\in\left(\frac{\tau\mu_{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})},\frac{\tau\mu^{*}}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}\right)\cap\left(0,\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)$.
@@
(3.15)${\geq}\Psi(y^{+})+\frac{1-\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2},$ (3.47)
@@
$V_{k-1}\geq\Psi(y^{k})+\frac{1-\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{2\alpha}{\left\|{y^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}.$ (3.48)
@@
When $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}=0$, $x$ and $z$ play no role in the algorithm. Similar to (3.47), we can obtain the following inequality:
$V_{\Lambda,\xi}(y,z,x)\geq\Psi(y^{+})+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$
Hence, if $\Psi$ has bounded level sets and $\beta<\frac{1}{\rho_{{}_{\rm{p}}}}$, we obtain the boundedness of $\{y^{k}\}$ and that ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$.
We will next show that accumulation points of the $x$ and $y$ sequence generated from Algorithm 1 are stationary points of $\Psi$. We say that a point $x^{*}$ is a stationary point of $\Psi$ (see (Wen et al., 2018, Definition 4.1)) if
$0\in\nabla f(x^{*})+\partial g(x^{*})+\nabla h(x^{*})+\partial p(x^{*}).$ (3.50)
Note that any local minimizer of $\Psi$ is a stationary point of $\Psi$ (see (Pham and Thi, 1997, Theorem 2(i))). To establish stationarity of accumulation points, we will prove that these points are fixed points of an operator associated with Algorithm 1, which we derive as follows. Observe that Algorithm 1 can be concisely written as fixed-point iterations of the operator $T_{\Lambda}:{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}$, given by
@@
where $T_{\Lambda}$ is defined as
@@
and use it as a measure of stationarity from the view that $R(y,z)=0$ if and only if $(y,z)\in\operatorname*{Fix}(T_{\Lambda})$. This residual function reduces to that in (Lee and Wright, 2019) when $f\equiv 0$ and $p\equiv 0$, and to that in (Liu and Takeda, 2022) when $f\equiv 0$. The next theorem shows that accumulation points of the $(y,z)$-sequence generated by Algorithm 1 are fixed points of $T_{\Lambda}$.
@@
Suppose that the hypotheses of Proposition 3.15 hold, and let $(x^{*},y^{*},z^{*})$ be an accumulation point of $\{(x^{k},y^{k},z^{k})\}$. Then $x^{*}=y^{*}$ and $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$.